solve using eliminaton
method
3x + 2y = 3
x - y = 6

Answer by checkley77(12844)

x-y=6 multiply this equation by 2 & add them.
3x+2y=3
2x-2y=12
----------------------
5x=15
x=15/5
x=3 answer.
3-y=6
-y=6-3
-y=3
y=-3 answer.
proof:
3*3+2*-3=3
9-6=3
3=3